OpenStudy (anonymous):

When solving the system x+2y=9 and x-y=6 which expression could be substituted for the x in the second equation? I think the answer is 9-x? is that correct?

OpenStudy (anonymous):

x = 9 -2y sub that in for x in the second eqn: (9-2y) -y =6 -3y = -3 y=1 Plug y=1 in to the first equation to get your x-value.

OpenStudy (anonymous):

x +2(1) =9 x = 7

OpenStudy (anonymous):

Does the method make sense to you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you